OpenStudy (anonymous):

Use the formula below to find the 5th term in the geometric sequence.

OpenStudy (anonymous):

\[a _{n}=(2)*(\frac{11}{10})^{n-1}\]

OpenStudy (anonymous):

A. \[\frac{ 14,641 }{ 5,000 }\] B. \[\frac{ 161,051 }{ 50,000 }\] C. \[\frac{ 176 }{ 5 }\] D. \[\frac{ 88 }{ 5 }\]